OpenStudy (anonymous):

Simplify 6^2. A. 12 B. 36 C. 62 -------------------- Simplify 10 ^3 -1 A. 99 B. 999 C. 729 Some help?

myininaya (myininaya):

6(6)=? 10^3=10(10)(10)=100(10)=1000 1000-1=?

OpenStudy (anonymous):

A For the first one

OpenStudy (anonymous):

B for the 2nd !

myininaya (myininaya):

A is wrong for the first one

OpenStudy (anonymous):

B

OpenStudy (anonymous):

Myiniaya i Major in Math Chemistry i can Redo it but i am Sure thats the answers

myininaya (myininaya):

Alex B is right for the first one since 6(6)=36

OpenStudy (anonymous):

6x6 Yes = 36 Oh well my Calculator on my home home one lied normally i triple check :P